OpenStudy (anonymous):

The size of a bicycle is determined by the diameter of the wheel. you want a 22 in. bicycle for your birthday. what is the area of one of the wheels ? A) 385.47 in. sq. B) 314.16 C) 484.37 D) 380.13

jimthompson5910 (jim_thompson5910):

Area = pi*radius^2 A = pi*r^2 Note: remember to cut the diameter in half to get the radius

OpenStudy (anonymous):

i dont know the radius cause they didnt even tell me t he diameter. @jim_thompson5910

jimthompson5910 (jim_thompson5910):

it states that "you want a 22 in. bicycle", so the diameter is 22 in

jimthompson5910 (jim_thompson5910):

since "The size of a bicycle is determined by the diameter of the wheel"

OpenStudy (anonymous):

oh okayy.

OpenStudy (anonymous):

i did 3.14*11 and got 34.54 then i squared that and didnt get any of the options available to me ?

jimthompson5910 (jim_thompson5910):

Remember that to follow the order of operations, you follow PEMDAS

jimthompson5910 (jim_thompson5910):

The E (for exponent) comes before M (multiplication), so you square before multiplying

jimthompson5910 (jim_thompson5910):

3.14*11^2 3.14*121 then you multiply

OpenStudy (anonymous):

i got 379.94. not an option..?

jimthompson5910 (jim_thompson5910):

hmm weird, I'm also getting 379.94

OpenStudy (anonymous):

it cant be wrong..

jimthompson5910 (jim_thompson5910):

hmm let me reread...one sec

OpenStudy (anonymous):

@SmoothMath helpp ?

jimthompson5910 (jim_thompson5910):

maybe if you used 22/7 for pi, it gets you somewhat closer

jimthompson5910 (jim_thompson5910):

oh wait, just type pi instead of 3.14 and you'll get a much better answer

jimthompson5910 (jim_thompson5910):

A = pi*r^2 A = pi*11^2 A = 121pi A = 380.13271108